Simplify 9-4[12-(-7+2)]
Answer by stanbon(75887)

Simplify 9-4[12-(-7+2)]
------------------
= 9 - 4[12-(-5)]
-----
= 9 - 4[17]
-----
= 9 - 68
-----
= -59
============
Cheers,
Stan H.
============